OpenStudy (anonymous):

how many solutions does this have? 4m-2n=16 and n=2m-3

OpenStudy (phi):

notice you can write n= 2m -3 as 3+n = 2m (add +3 to both sides) 3= 2m - n (add -n to both sides) multiply both sides by 2 6 = 2(2m-n) 6 = 4m -2n or 4m - 2n = 6 now compare this to the other equation 4m-2n= 16

OpenStudy (anonymous):

ok what am i sopossed to see

OpenStudy (phi):

the first equation says, there are two numbers m and n, and when you multiply m by 4 and subtract 2 times n you get 6 the second equation says, "do the same thing and you get 16" in other words, both equations cannot be true at the same time. a solution to two equations is something that makes both true *at the same time*

OpenStudy (anonymous):

so theres only one solution?

OpenStudy (phi):

Here is another way to do this 4m-2n=16 and n=2m-3 replace n with 2m-3 (because the 2nd equation says n is the same as 2m-3) 4m - 2(2m-3)= 16 distribute the -2: 4m -4m + 6 =16 4m - 4m is 0, so this simplifies to 6 = 16 which is not true. That is a sign there is no solution

OpenStudy (anonymous):

o ok this helps a lot thanks for the help